Solve the system of equations. (If the system is dependent, enter a general solution in terms of c. If there is no solution
{x+5y− 4z= 22
4x − y+ z= 3
5x+ 4y −5z=35
(x, y, z) =

Answer by Alan3354(69443)

Solve the system of equations. (If the system is dependent, enter a general solution in terms of c.

x  + 5y - 4z = 22
4x  - y +  z = 3
5x + 4y - 5z = 35
-------
Add the 1st and 2nd eqns.
 x + 5y - 4z = 22
4x  - y +  z = 3
-----------------------  Add
5x + 4y - 3z = 25
5x + 4y - 5z = 35  ----  eqn 3
-------------------------------------- Subtract
2z = -10
z = -5
===========
Pick any 2 of the equations.
x  + 5y - 4z = 22
4x  - y +  z = 3
-------
z = -5
x  + 5y + 20 = 22
4x  - y + -5 = 3
---
x  + 5y = 2 times 4 --> 4x + 20y = 8
4x  - y = 8
---------
4x + 20y = 8
4x - y = 8
---------------- Subtract
21y = 0
y = 0
x = 2
---> (2,0,-5)